# Time-dependent flux backgrounds in type IIB supergravity

### v1

@@ 1 Introduction and summary
The type IIB supergravity appears in the low energy limit of the type IIB superstring theory in ten dimensions Polchinski:1998rr . Its bosonic sector consists of the 10D metric field, a complex zero-form axiodilaton field, along with a self-dual five-form and a complex three-form field strengths. Finding its classical backgrounds amounts to solving the 10D Einstein field equations and the equations of motion and Bianchi identities for the form fields, collectively referred to as the type IIB supergravity equations. Over the years various classes of such backgrounds have been constructed, including solutions with an $AdS_{n}$ component (see e.g. Macpherson:2014eza ; Figueroa-OFarrill:2012whx ), black hole solutions (see e.g. Duff:1999rk ; Mandal:2000rp ; Emparan:2006mm ), pp-wave solutions (see e.g. Blau:2001ne ; Blau:2002dy ; Bena:2002kq ; Gran:2018ijr ), etc. Our interests are classical backgrounds that allow for a noncompact time-dependent 4D spacetime component. In this regard there are no-go theorems deWit:1986mwo ; Maldacena:2000mw ; Gibbons:2003gb ; HariDass:2002si (see Russo:2019fnk ; Das:2019vnx ; Bernardo:2022ony ; Faruk:2024usu ; Shukla:2019dqd ; Niedermann:2025tuo for recent explorations) concerning the existence of a 4D component with nonnegative Ricci scalar, such as Minkowski, Kasner or de Sitter spacetimes, based on specific ansatzes for the metric and other field content. [^1] Allowing local sources, especially objects with negative tension, can loosen such no-go theorems and allow for a 4D spacetime component with zero Ricci scalar such as Minkowski or Kasner spacetimes Giddings:2001yu . Notably, some of the above discussions on no-go theorems employ various energy conditions; for review on the consequences and known violations of such energy conditions, see Kontou:2020bta ; Curiel:2014zba ; Dain:2013vsa ; Malament:2005rc .
@@
In general, a systematic tabulation of time-dependent 4D components allowed in supergravity and string theory, together with an analysis of their cosmological implications, will be of considerable value. Attempts have been made to construct such backgrounds, either in toy supergravity models or by switching off all fluxes Chen:2002yq ; Bhattacharya:2003sh ; Neupane:2009jn ; Mueller:1989in ; Townsend:2003fx ; Ohta:2003uw ; Ohta:2003pu ; Roy:2003nd ; Roy:2006du ; Roy:2007fu ; Nayek:2014iza ; Nayek:2017yqx ; Sabra:2020gio ; Fakhredin:2025ngo . In these discussions, models of higher-dimensional gravity coupled to a dilaton scalar field and a higher-form potential field have been considered, thus incorporating some ingredients found in various supergravity theories. However, they miss, for example, Chern-Simons terms, as well as the full intricacy of the supergravity equations arising from multiple form fields (which couple to the dilaton differently). See Buchel:2002kj ; SenGupta:2007lbr for time-dependent examples constructed using different methods, such as gauge/gravity correspondence and analytical continuation of known static solutions. While our focus in this paper is on type IIB supergravity, see Billyard:1999dg ; Cavaglia:2000rx ; Cavaglia:2000uu ; Cavaglia:2001sb for a discussion of 4D cosmologies ignoring only the R–R one-form potential in 10D type IIA supergravity, while for more general constructions see Marconnet:2022fmx . [^3] Type IIA supergravity, which lacks self-dual field strengths, arises from compactification of 11D M-theory on an internal circle.
@@
  • In our analysis we keep a time-dependent scale factor multiplying the 6D internal space ($\tilde{g}_{mn}$), and field strengths with components only partially covering the 4D noncompact spacetime ($\hat{g}_{\mu\nu}$). We identify the 4D Einstein frame through a Weyl rescaling of the 4D metric, which involves the aforementioned time-dependent scale factor. Notably, we do not consider any local sources in constructing backgrounds.
@@
  • More generally, our constructions are 4D Bianchi type I universes with negative Ricci scalar; however, special cases yield spatially flat FLRW universes. In some examples, the 4D scale factors exhibit power-law growth at late times. We also discuss the supergravity constraints on the 6D scale factor.
@@
  • Next, we revisit the Maldacena–Nuñez no-go analysis, incorporating new elements that appear in our constructions, namely an overall noncompact spacetime-dependent scale factor multiplying the internal metric, and field strengths with components partially covering the noncompact directions. For this analysis (which is largely independent from above), we also include a warp factor dependent on the internal coordinates, which multiplies the noncompact spacetime. We argue that, with these generalizations, a four-dimensional Einstein-frame metric with positive Ricci scalar cannot be ruled out solely by such an analysis.
@@ 2.2.3 $\tilde{F}_{5}=0,\ G_{3},\tau\neq 0,\ \tau(t)$
We now turn to the $mn$-components of the Einstein field equations. If we demand the internal manifold to be Ricci flat, i.e., $\tilde{R}_{mn}=0$, we have $\eta_{m}{}^{\tilde{p}}\bar{\eta}_{np}\propto\tilde{g}_{mn}$, which follows from the factorization of the $x$-dependence from the $y$-dependence in (7). Consistent with the harmonicity of $\eta_{2}$, on a compact Kähler internal manifold we can simply take [^11] See footnote 8.
@@ 2.2.4 $\textrm{Self-dual}\ \tilde{F}_{5},\ G_{3},\tau\neq 0,\ \tau(t)$
We now turn to the $mn$-components of the Einstein field equations. If we demand the internal manifold to be Ricci flat, i.e., $\tilde{R}_{mn}=0$, we have $\lambda_{m}{}^{\tilde{p}}\lambda_{np}\propto\tilde{g}_{mn}$, which follows from the factorization of the $x$-dependence from the $y$-dependence in (7). Consistent with the harmonicity of $\lambda_{2}$, on a compact Kähler internal manifold we can simply take [^13] See footnote 8.
@@ 3.2 Checks in 4D
written generally; however, in our explicit constructions, not all vectors from $\{\gamma_{\mu},\zeta_{\mu},\vartheta_{\mu},\partial_{\mu}\tau\}$ contribute simultaneously within a single case. Note that a hat on raised indices refers to the contraction performed using $\hat{g}^{E}_{\mu\nu}$, which is a convention restricted to this subsection and applies to all subsequent equations. [^34] Aside from the transformation formula between 4D frames given in footnote 33.
@@ 4 Revisiting Maldacena–Nuñez analysis
$=\frac{1}{2\kappa_{10}^{2}}\int d^{6}y\sqrt{\tilde{g}}\ e^{2A}\int d^{4}x\sqrt{-\hat{g}^{E}}\ e^{\Omega}e^{6\beta}\hat{R}^{E}+\cdots\,,\quad\hat{g}_{\mu\nu}=e^{\Omega(x)}\hat{g}^{E}_{\mu\nu}\,.$
@@
where $\kappa_{4}$ is a constant parameter. In this case, the 4D gravitational constant depends on the warp factor $A(y)$.
@@ 2. SUGRA constraints on the scale factor multiplying the internal metric:
which does not necessarily approach zero at late times, as the constant $k>4/a$ can be chosen to avoid this, given the constant $a>0$ that appears in the relevant constructions. We now provide a proof that the factors above in $e^{6\beta+\sigma}$ do not affect the statement that when $\beta(t)$ approaches to a constant asymptotically, $e^{6\beta+\sigma}$ approaches zero – a statement we used to discard such functions. For such a $\beta(t)$, the late-time behaviour is generically of the form: $\beta\sim\beta_{0}+\beta_{n}t^{-n}+\mathcal{O}(t^{n+1})$ for some $n>0$ (even fractions are allowed) and constants $\beta_{0},\beta_{n}$. Moreover, we should have $e^{b\beta_{0}}<b^{2}$, although this does not affect the leading power of $t$ in $e^{6\beta+\sigma}$ at late times. Based on the above formula, we obtain $e^{6\beta+\sigma}\sim t^{-2(n+1)}$ (up to a positive proportionality constant).
@@
which does not necessarily approach zero at late times, as the constant $k>2$ can be chosen to avoid this in the relevant construction. Similar to the previous case, at late times $e^{6\beta+\sigma}\sim t^{-2(n+1)}$ when $\beta\sim\beta_{0}+\beta_{n}t^{-n}+\mathcal{O}(t^{n+1})$ for some $n>0$.
@@ 4. Homogeneous, anisotropic, and FLRW cosmologies
Below (37), we have discussed the spatial scale factors $e^{6\beta(t(t^{\prime}))+\alpha_{i}(t(t^{\prime}))}$ of the 4D Einstein frame metric in cosmic time $t^{\prime}$. In our constructions, generically the 4D Einstein frame metric is a Bianchi type I universe, i.e., spatially homogeneous yet anisotropic (for references, see ToledoSesma:2015wxj ; Verma:2024lex ). However, for flux configurations A and B, our constructions allow all the spatial scale factor exponents $\alpha_{i}$ to be equal, so the Einstein frame metric takes a spatially flat FLRW form in those cases, as argued in section 2.3 along with providing the corresponding choices of constant parameters. In contrast, for flux configuration C, only two of the spatial scale factors can be made equal. In all our constructions the Ricci scalar of the 4D Einstein frame metric is negative, $\hat{R}^{E}<0$. Due to this, our backgrounds could be viable candidates for a kination epoch in the early universe. The latter also requires moduli fields in the 4D EFT whose kinetic energy dominates the energy density. Recently, Apers:2024dtn ; Apers:2022cyl studied perturbations of 10D Kasner backgrounds, relating them to rolling moduli in the 4D kination epoch by compactifying on a 6D torus. A study of the 4D EFT of the geometric moduli of the 6D internal metric of our constructed backgrounds could therefore be relevant in this context.
@@ 5. Time-dependent warped flux backgrounds
Here, by warped we refer to a nonzero $A(y)$ as in (73), while we continue to refer to $\beta(x)$ as the 6D scale factor exponent. In this work, all the constructions work with $A(y)=0$. However, our quest is to construct viable time-dependent 4D spacetimes within the framework of 10D supergravity theory, regardless of other ingredients like $A(y)$. One time-dependent background with $A(y)\neq 0$, but $\beta(x)=0$, arises in the GKP setup Giddings:2001yu . As also pointed out in Frey:2024jqy , any 4D Ricci-flat metric (with conformally Ricci-flat internal manifold) solves type IIB supergravity equations in presence of D-branes and O-planes, and self-dual $\tilde{F}_{5}$ featuring $[4,1]$ and $[0,5]$ components and imaginary self-dual $G_{3}$ featuring harmonic $[0,3]$ components. This can be checked straightforwardly. Thus, for example, a 4D Kasner metric is a viable time-dependent 4D metric as it is Ricci-flat. Note that the GKP fluxes are of MN-type, which we did not consider in the current work. It would be important to solve the supergravity equations while keeping both $A(y),\beta(x)$ nontrivial, allowing both MN and non-MN type fluxes along with local sources, and searching for allowed time-dependent 4D spacetimes. We leave such a detailed scan for future work. Obtaining a time-dependent background that approaches a GKP background with a warped 4D Minkowski component at late times will be important, as such backgrounds are of particular interest in particle physics phenomenology (for reviews, see Agrawal:2022rqd ; Maharana:2012tu ; Cicoli:2023opf ).
@@ 6. Positive Ricci scalar of 4D Einstein frame metric
In absence of the 6D scale factor, non-MN types fluxes, and O-planes (which are objects with negative tension), (78) rules out $\hat{R}^{E}\geq 0$ generically across any supergravity theories. [^45] Note that in absence of the 6D scale factor, $\hat{R}^{E}=\hat{R}$. GKP Giddings:2001yu by inclusion of O-planes constructs 4D spacetimes with $\hat{R}^{E}=0$ in type IIB supergravity. Our generalizations of the Maldacena–Nuñez analysis incorporating the 6D scale factor, non-MN types fluxes, do not rule out the existence of a positive $\hat{R}^{E}$ solely by considering the traced Einstein equations (see, (78) and (80)). There are two approaches one can take to attempt drawing conclusive remarks: (i) by imposing energy conditions on top of the traced Einstein equations, along the lines of Russo:2019fnk ; Das:2019vnx ; Bernardo:2021zxo ; Bernardo:2022ony ; Faruk:2024usu , or (ii) by conducting a comprehensive search in full generality, in which all components of the field strengths and both warp factors $A(y)$ and $\beta(x)$ are retained, and exploring all supergravity theories. However, note that energy conditions are not absolute principles; there are known violations of all pointwise energy conditions by physically reasonable fields Kontou:2020bta ; Curiel:2014zba .

### v2

@@ 1 Introduction and summary
The type IIB supergravity appears in the low energy limit of the type IIB superstring theory in ten dimensions Polchinski:1998rr . Its bosonic sector consists of the 10D metric field, a complex zero-form axiodilaton field, along with a self-dual five-form and a complex three-form field strengths. Finding its classical backgrounds amounts to solving the 10D Einstein field equations and the equations of motion and Bianchi identities for the form fields, collectively referred to as the type IIB supergravity equations. Over the years various classes of such backgrounds have been constructed, including solutions with an $AdS_{n}$ component (see e.g. Macpherson:2014eza ; Figueroa-OFarrill:2012whx ), black hole solutions (see e.g. Duff:1999rk ; Mandal:2000rp ; Emparan:2006mm ), pp-wave solutions (see e.g. Blau:2001ne ; Blau:2002dy ; Bena:2002kq ; Gran:2018ijr ), etc. Our interests are classical backgrounds that allow for a noncompact time-dependent 4D spacetime component. In this regard there are no-go theorems deWit:1986mwo ; Maldacena:2000mw ; Gibbons:2003gb ; HariDass:2002si (see Russo:2019fnk ; Das:2019vnx ; Bernardo:2022ony ; Faruk:2024usu ; Shukla:2019dqd ; Niedermann:2025tuo for recent explorations) concerning the existence of a 4D component with nonnegative Ricci scalar, such as Minkowski, Kasner or de Sitter spacetimes, based on specific ansatzes for the metric and other field content. [^1] Allowing local sources, especially objects with negative tension, can loosen such no-go theorems and allow for a 4D spacetime component with zero Ricci scalar such as Minkowski or Kasner spacetimes Giddings:2001yu . In the presence of such branes, whether a 4D de Sitter spacetime can be realized has been discussed in Dasgupta:2014pma , adopting similar types of 10D metric and fluxes. Notably, some of the above discussions on no-go theorems employ various energy conditions; for review on the consequences and known violations of such energy conditions, see Kontou:2020bta ; Curiel:2014zba ; Dain:2013vsa ; Malament:2005rc .
@@
In general, a systematic tabulation of time-dependent 4D components allowed in supergravity and string theory, together with an analysis of their cosmological implications, will be of considerable value. Attempts have been made to construct such backgrounds, either in toy supergravity models or by switching off all fluxes Chen:2002yq ; Bhattacharya:2003sh ; Neupane:2009jn ; Mueller:1989in ; Townsend:2003fx ; Ohta:2003uw ; Ohta:2003pu ; Roy:2003nd ; Roy:2006du ; Roy:2007fu ; Nayek:2014iza ; Nayek:2017yqx ; Sabra:2020gio ; Fakhredin:2025ngo . In these discussions, models of higher-dimensional gravity coupled to a dilaton scalar field and a higher-form potential field have been considered, thus incorporating some ingredients found in various supergravity theories. However, they do not have, for example, contributions from Chern-Simons terms, as well as the full intricacy of the supergravity equations arising from multiple form fields (which couple to the dilaton differently). See Buchel:2002kj ; SenGupta:2007lbr for time-dependent examples constructed using different methods, such as gauge/gravity correspondence and analytical continuation of known static solutions. While our focus in this paper is on type IIB supergravity, see Billyard:1999dg ; Cavaglia:2000rx ; Cavaglia:2000uu ; Cavaglia:2001sb for a discussion of 4D cosmologies ignoring only the R–R one-form potential in 10D type IIA supergravity, while for more general constructions see Marconnet:2022fmx . [^3] Type IIA supergravity, which lacks self-dual field strengths, arises from compactification of 11D M-theory on an internal circle. Moreover, cosmological solutions in supergravity systems coupled to brane sources have been constructed previously (see, e.g., Chen:2005ae ), while the present work does not incorporate branes.
@@
  • In our analysis we keep a time-dependent scale factor multiplying the 6D internal space ($\tilde{g}_{mn}$), and field strengths with components only partially covering the 4D noncompact spacetime ($\hat{g}_{\mu\nu}$) – such fluxes are referred to as non-MN type. [^4] The fluxes used are time-dependent and not purely internal, and thus standard flux quantization conditions do not apply. We identify the 4D Einstein frame through a Weyl rescaling of the 4D metric, which involves the aforementioned time-dependent scale factor. Notably, we do not consider any local sources in constructing backgrounds.
@@
  • More generally, our constructions are 4D Bianchi type I universes with negative Ricci scalar; however, special cases yield spatially flat FLRW universes (thereby demonstrating that the isometries of FLRW spacetime can be retained even in the presence of non-MN type fluxes). In some examples, the 4D scale factors exhibit power-law growth at late times.
@@
  • Next, we revisit the Maldacena–Nuñez no-go analysis, incorporating new elements that appear in our constructions, namely an overall noncompact spacetime-dependent scale factor multiplying the internal metric, and field strengths with components partially covering the noncompact directions. For this analysis (which is largely independent from above), we also include a warp factor dependent on the internal coordinates, which multiplies the noncompact spacetime. We argue that, with these generalizations, a 4D Einstein-frame metric with positive Ricci scalar cannot be ruled out solely by such an analysis. In this regard, even the contribution of the 6D scale factor is individually notable.
@@ 2.2.3 $\tilde{F}_{5}=0,\ G_{3},\tau\neq 0,\ \tau(t)$
We now turn to the $mn$-components of the Einstein field equations. If we demand the internal manifold to be Ricci flat, i.e., $\tilde{R}_{mn}=0$, we have $\eta_{m}{}^{\tilde{p}}\bar{\eta}_{np}\propto\tilde{g}_{mn}$, which follows from the factorization of the $x$-dependence from the $y$-dependence in (7). Consistent with the harmonicity of $\eta_{2}$, on a compact Kähler internal manifold we can simply take [^12] See footnote 9.
@@ 2.2.4 $\textrm{Self-dual}\ \tilde{F}_{5},\ G_{3},\tau\neq 0,\ \tau(t)$
We now turn to the $mn$-components of the Einstein field equations. If we demand the internal manifold to be Ricci flat, i.e., $\tilde{R}_{mn}=0$, we have $\lambda_{m}{}^{\tilde{p}}\lambda_{np}\propto\tilde{g}_{mn}$, which follows from the factorization of the $x$-dependence from the $y$-dependence in (7). Consistent with the harmonicity of $\lambda_{2}$, on a compact Kähler internal manifold we can simply take [^14] See footnote 9.
@@ 3.2 Checks in 4D
written generally; however, in our explicit constructions, not all vectors from $\{\gamma_{\mu},\zeta_{\mu},\vartheta_{\mu},\partial_{\mu}\tau\}$ contribute simultaneously within a single case. Note that a hat on raised indices refers to the contraction performed using $\hat{g}^{E}_{\mu\nu}$, which is a convention restricted to this subsection and applies to all subsequent equations. [^35] Aside from the transformation formula between 4D frames given in footnote 34.
@@ 4 Revisiting Maldacena–Nuñez analysis
$=\frac{1}{2\kappa_{10}^{2}}\int d^{6}y\sqrt{\tilde{g}}\ e^{2A}\int d^{4}x\sqrt{-\hat{g}^{E}}\ e^{\Omega}e^{6\beta}\hat{R}^{E}+\cdots\,,\quad\hat{g}_{\mu\nu}=e^{\Omega(x)}\hat{g}^{E}_{\mu\nu}\,.$ (76)
@@
where $\kappa_{4}$ is a constant parameter. In this case, the 4D gravitational constant depends on the warp factor A(y)A(y). [^40] This is given by, $(16\pi G_{N})^{-1}=\left(e^{-2\kappa_{4}}/2\kappa_{10}^{2}\right)\int d^{6}y\sqrt{\tilde{g}}\ e^{2A}$.
@@ 2. SUGRA constraints on the scale factor multiplying the internal metric:
which does not necessarily approach zero at late times, as the constant $k>4/a$ can be chosen to avoid this, given the constant $a>0$ that appears in the relevant constructions. We now provide a proof that the factors above in $e^{6\beta+\sigma}$ do not affect the statement that when $\beta(t)$ approaches to a constant asymptotically, $e^{6\beta+\sigma}$ approaches zero – a statement we used to discard such functions. For such a $\beta(t)$, the late-time behaviour is generically of the form: $\beta\sim\beta_{0}+\beta_{n}t^{-n}+\mathcal{O}(t^{-n-1})$ for some $n>0$ (even fractions are allowed) and constants $\beta_{0},\beta_{n}$. Moreover, we should have $e^{b\beta_{0}}<b^{2}$, although this does not affect the leading power of $t$ in $e^{6\beta+\sigma}$ at late times. Based on the above formula, we obtain $e^{6\beta+\sigma}\sim t^{-2(n+1)}$ (up to a positive proportionality constant).
@@
which does not necessarily approach zero at late times, as the constant $k>2$ can be chosen to avoid this in the relevant construction. Similar to the previous case, at late times $e^{6\beta+\sigma}\sim t^{-2(n+1)}$ when $\beta\sim\beta_{0}+\beta_{n}t^{-n}+\mathcal{O}(t^{-n-1})$ for some $n>0$.
@@ 3. Late time 4D scale factors and Kasner-like backgrounds
For the cases with power-law growth of the temporal scale factor $e^{6\beta+\sigma}$ in $t$, the integral defining the cosmic time $t^{\prime}(t)$ grows as a power of $t$ (with the power raised by one) at large $t$. While, for the cases with exponential growth of the temporal scale factor in $t$, $t^{\prime}(t)$ grows exponentially with $t$ at large $t$. Therefore, at late cosmic time $t^{\prime}$ (which corresponds to large $t$, as discussed previously), the inversion $t(t^{\prime})$ grows as a power of $t^{\prime}$ or logarithmically in $t^{\prime}$, respectively in these two types of cases. Consequently, in the second type of cases, the spatial scale factors that grow exponentially in $t$ also evolve as power laws in cosmic time $t^{\prime}$, as in the first type. It will be interesting to study the spatial scale factors $e^{6\beta(t(t^{\prime}))+\alpha_{i}(t(t^{\prime}))}$ (e.g., those appear in appendix D) in cosmic time $t^{\prime}$ by numerically inverting the integral that defines $t^{\prime}(t)$.
@@ 4. Homogeneous, anisotropic, and FLRW cosmologies
Below (37), we have discussed the spatial scale factors $e^{6\beta(t(t^{\prime}))+\alpha_{i}(t(t^{\prime}))}$ of the 4D Einstein frame metric in cosmic time $t^{\prime}$. In our constructions, generically the 4D Einstein frame metric is a Bianchi type I universe, i.e., spatially homogeneous yet anisotropic (for references, see ToledoSesma:2015wxj ; Verma:2024lex ). However, for flux configurations A and B, our constructions allow all the spatial scale factor exponents $\alpha_{i}$ to be equal, so the Einstein frame metric takes a spatially flat FLRW form in those cases, as argued in section 2.3 along with providing the corresponding choices of constant parameters. In contrast, for flux configuration C, only two of the spatial scale factors can be made equal. In all our constructions the Ricci scalar of the 4D Einstein frame metric is negative, $\hat{R}^{E}<0$. [^47] In the cases we obtain a 4D FLRW universe ($\alpha_{i}=\alpha$), one can check that the second derivative of the 3D spatial scale factor w.r.t. the cosmic time $t^{\prime}$, computed as $e^{-3\beta-\sigma/2}\frac{d}{dt}[e^{-3\beta-\sigma/2}\frac{d}{dt}e^{3\beta+\alpha/2}]$, is negative. Due to this, our backgrounds could be viable candidates for a kination epoch in the early universe. The latter also requires moduli fields in the 4D EFT whose kinetic energy dominates the energy density. Recently, Apers:2024dtn ; Apers:2022cyl studied perturbations of 10D Kasner backgrounds, relating them to rolling moduli in the 4D kination epoch by compactifying on a 6D torus. A study of the 4D EFT of the geometric moduli of the 6D internal metric of our constructed backgrounds could therefore be relevant in this context.
@@ 5. Time-dependent warped flux backgrounds
Here, by warped we refer to a nonzero $A(y)$ as in (73), while we continue to refer to $\beta(x)$ as the 6D scale factor exponent. In this work, all the constructions work with $A(y)=0$. However, our quest is to construct viable time-dependent 4D spacetimes within the framework of 10D supergravity theory, regardless of other ingredients like $A(y)$. One time-dependent background with $A(y)\neq 0$, but $\beta(x)=0$, arises in the GKP setup Giddings:2001yu . As also pointed out in Frey:2024jqy , any 4D Ricci-flat metric (with conformally Ricci-flat internal manifold) solves type IIB supergravity equations in presence of D-branes and O-planes, and self-dual $\tilde{F}_{5}$ featuring $[4,1]$ and $[0,5]$ components and imaginary self-dual $G_{3}$ featuring harmonic $[0,3]$ components. This can be checked straightforwardly. Thus, for example, a 4D Kasner metric is a viable time-dependent 4D metric as it is Ricci-flat. Note that the GKP fluxes are of MN-type, which we did not consider in the current work. It would be important to solve the supergravity equations while keeping both $A(y),\beta(x)$ nontrivial, allowing both MN and non-MN type fluxes along with local sources, and searching for allowed time-dependent 4D spacetimes. We leave such a detailed scan for future work. Obtaining a time-dependent background that approaches a GKP background with a warped 4D Minkowski component at late times will be important, as such backgrounds are of particular interest in particle physics phenomenology (for reviews, see Agrawal:2022rqd ; Maharana:2012tu ; Cicoli:2023opf ). Moreover, effects of quantum corrections including higher-order curvature corrections (e.g., Dasgupta:2014pma ; Dasgupta:2018rtp ; Dasgupta:2019vjn ) on the allowed 4D spacetimes $\hat{g}^{E}_{\mu\nu}(t)$, regardless of the sign of $\hat{R}^{E}(t)$, will be important to study for early universe cosmologies.
@@ 6. Positive Ricci scalar of 4D Einstein frame metric
In absence of the 6D scale factor, non-MN types fluxes, and O-planes (which are objects with negative tension), (79) rules out $\hat{R}^{E}\geq 0$ generically across any supergravity theories. [^48] Note that in absence of the 6D scale factor, $\hat{R}^{E}=\hat{R}$. GKP Giddings:2001yu by inclusion of O-planes constructs 4D spacetimes with $\hat{R}^{E}=0$ in type IIB supergravity. Our generalizations of the Maldacena–Nuñez analysis incorporating the 6D scale factor, non-MN types fluxes, do not rule out the existence of a positive $\hat{R}^{E}$ solely by considering the traced Einstein equations (see, (79) and (81)). Even the contribution of the 6D scale factor to $\hat{R}^{E}$ is noteworthy by itself. There are two approaches one can take to attempt drawing conclusive remarks: (i) by imposing energy conditions on top of the traced Einstein equations, along the lines of Russo:2019fnk ; Das:2019vnx ; Bernardo:2021zxo ; Bernardo:2022ony ; Faruk:2024usu , or (ii) by conducting a comprehensive search in full generality, in which all components of the field strengths and both warp factors $A(y)$ and $\beta(x)$ are retained, and exploring all supergravity theories (even with branes included, as (79), (81) hold for any contributions to the energy-momentum tensor). However, note that energy conditions are not absolute principles; there are known violations of all pointwise energy conditions by physically reasonable fields Kontou:2020bta ; Curiel:2014zba ; Visser:1999de ; Barcelo:1999hq ; Barcelo:2000zf .
Notably, solution with a 4D component with $\hat{R}^{E}>0$ to higher dimensional vacuum Einstein equations (i.e., with zero fluxes) was reported earlier, considering non-trivial time-dependent $\beta(t)$ and $A(y)=0$, however a compact hyperbolic internal space was taken Townsend:2003fx . A negative scalar curvature of the internal space contributes positively to $\hat{R}^{E}$, see (125). Our (79) does not directly see $\tilde{R}$, as in original Maldacena–Nuñez analysis. However, (81) does see it and feeds back into $\beta(x)$, which in turn affects (79). It will be interesting to study if hyperbolic internal spaces are necessary in constructions with $\hat{R}^{E}>0$ in presence of fluxes.